OpenStudy (anonymous):

determine the surface area of the surface given parametrically by r(u,v) = , 0 is less than or equal to u which is less than or equal to 1, and 0 is less than or equal v which is less than or equal to 1.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

i surface integral is similar to a line integral ....murmur

OpenStudy (amistre64):

i spose it begins as the sqrt([dx]^2 + [dy]^2 + [dz]^2) right?

OpenStudy (anonymous):

it's a step somewhere.

OpenStudy (amistre64):

wish i could do more; but thats as far as I can get in this one :)

OpenStudy (amistre64):

at least the bounds are constants :)

OpenStudy (amistre64):

x= u+v y = v z = u x = z+y maybe? x' = z' + y'

OpenStudy (amistre64):

your gonna have to hope a smarter then me comes along ;)

OpenStudy (amistre64):

or at least chimes in ;)

OpenStudy (anonymous):

it's no problem. thanks for trying tho. there is a question you might can answer for me. when using a determinant to find the cross product of 2 vectors, which way do you cross multiply first? bottom left to top right or bottom right to top left?

OpenStudy (amistre64):

its just the nature of the best; the matrix itself is just a way to organize the information; but the results of acutally going thru the work results in the technique

OpenStudy (amistre64):

the crossing of vectors mimics the determinate of a matrix; so its a good visual to follow

OpenStudy (amistre64):

but spose we have 2 vectors on a plane froma single point and we want to determine the normal to the plane... we would set up the proper dots products to get a zero from each vector and the resulting summations result in a form tha tmatches the determinate of a matrix...

OpenStudy (amistre64):

its alot to try to type in here :) but if you need more clarification, just let me know

OpenStudy (anonymous):

I think it's \[\int\int | {\delta r \over \delta x} \times {\delta r \over \delta y}|dxdy\] Where r is what you have defined there parametrically (so you'll probably need a Jacobian as well).

OpenStudy (anonymous):

isn't the cross product the jacobian?

OpenStudy (anonymous):

your equation is correct according to what i just found.

OpenStudy (anonymous):

Oh.. I'm sorta right.. This should help some: http://www.math24.net/surface-integrals-of-first-kind.html

OpenStudy (anonymous):

thanks for the link. i found the answer i needed. I found the definition of the surface area.

OpenStudy (anonymous):

In particular this seems promising: \[\int\int_SdS = \int\int_{D(u,v)}r(u,v)| {\delta r \over \delta u} \times {\delta r \over \delta v}|dudv\]

OpenStudy (anonymous):

that's what i found. thanks for your help.